REU 2006: REVISED SCHEDULE FOR WEEK TWO

ALL TALKS WILL BE IN RYERSON 251 EXCEPT MWF AFTERNOONS IN ECKHART 133

SECOND WEEK:

NOTE: attendance at the YSP and SESAME morning preparation and training sessions by all people who will be counsellors in either of these programs is required.

Date  YSP/SES Abert/Babai  Abert  Babai  Farb  May
Apprentice Problems Discrete Geometry K-theory
Mon, June 26 9:00–12:00 9:30–12:00 1:00–3:00 3:30–5:00
Tue, June 27 9:00–12:00 9:30–12:00 3:30–5:00  1:00–3:00
Wed, June 28 9:00–12:00 9:30–12:00 1:00–3:0 3:30–5:00
Thu, June 29 9:00–12:00 9:30–12:00 3:30–5:00  1:00–3:00
Fri, June 30  9:00–12:00 9:30–12:00  1:00–3:00  3:30–5:00

TUTORIALS: 5:00 onwards

Here is the schedule in summary form:

SECOND WEEK:

YSP/SESAME Preparation: 9:00–12:00 M-F Abert/Babai (Apprentice): 9:30–12:00 M-F Abert (Problems): 3:30–5:00 TTh Babai (Discrete mathematics): 1:00–3:00 MWF Farb (Geometry and topology): 1:00–3:00 TTh May (K-theory): 3:30-5:00 MWF